# DIO: Refining Mutual Information and Causal Chain to Enhance Machine Abstract Reasoning Ability

### v4

@@ VIII-A Target Reasoning Rather Than Solving Problems
This paper realizes that setting maximization of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective of DIO is suboptimal. Although this statistical measure strongly correlates with the solving accuracy of RPM tasks, it neglects the human-predefined reasoning logic [35], yielding a model optimized for high accuracy rather than for faithful causal reasoning. Thus, the performance gains brought by mutual information may be superficial. Both the Brando and WORLD methodologies aim to assist DIO in maximizing mutual information, yet neither resolves this limitation. In a nutshell, mutual-information-driven approaches are effective but need an explicit causal perspective.
@@
In order to verify semantic alignment, this paper takes the “Distribute 9” dataset from the RPM-style RAVEN benchmark [18] as the research object. We conduct training on this dataset using DIO, as well as DIO supplemented with Brando and WORLD. Subsequently, we extract 30 instances that share strictly identical progressive patterns and compute the pairwise similarities among the progressive pattern representations that DIO encodes for these 30 instances. The calculated $30\times 30$ similarity matrices are visualized as heatmaps in Figure 9.
Figure 9: Heatmap of Progressive Pattern Representation Extracted from the Distribute 9 under RAVEN Problems.
@@
In preliminary experiments, we found that DIO exhibits deviations in modeling the segment of the causal chain that involves “abstract attributes $\rightarrow$ progressive patterns of attributes”. This corresponds to a misalignment between the outputs $P_{\alpha{jkl}}$ produced by Formula (IV-C) and human semantics. Therefore, this paper proposes rectifying the progressive pattern representations $P_{\alpha{jkl}}$ encoded by DIO as an effective remedy.
@@
To this end, we propose the DIEGO method. DIEGO constructs a set of trainable reference vectors $\{V_{f}|f\in[1,F]\}$. By assigning each vector a distinct metadata modality, this entire set can be considered as a comprehensive metadata dictionary. In addition, DIEGO introduces an auxiliary loss term $\ell_{\text{DIEGO}}$ that serves to align each progressive pattern representation $P_{\alpha{jkl}}$ extracted by the DIO with its corresponding reference vector in $\{V_{f}|f\in[1,F]\}$ based on the accompanying metadata. This term is formulated as follows:
@@
Here, $\text{Meta}_{\hat{f}l}$ is a binary indicator that flags whether the $l$-th metadata entry matches the modality associated with the $\hat{f}-$th vector within the reference dictionary $\{V_{f}|f\in[1,F]\}$. Meanwhile $\tau$ is a trainable scaling coefficient. The detailed formulation of $\ell_{\text{DIEGO}}$ elucidates our alignment mechanism. Guided by the metadata-derived indicators, this loss term employs logarithmic smoothing and exponential scaling to regulate the cosine similarities between each $P_{\alpha_{jkl}}$ and all vectors in the reference dictionary. Evidently, our objective has shifted from statistical quantity control to vector similarity control, distinguishing DIEGO from Brando and WORLD: it is no longer mutual-information-driven but progressive-pattern-driven. When the DIO model is trained with the composite loss $\ell_{\text{DIO}}+\ell_{\text{DIEGO}}$ and the trainable reference dictionary $\{V_{f}|f\in[1,F]\}$, the forward pass of the relevant modules proceeds as illustrated in Figure 10.
@@
It is widely accepted that introducing more supervisory signals should improve model performance. Indeed, previous studies have attempted to leverage metadata as additional supervision, aiming to boost both interpretability and reasoning accuracy simultaneously. Recent work [23, 26, 25, 29] has collectively shown, however, that this practice can backfire: directly embedding metadata into proposed models actually degrades reasoning performance. This paper argues that the DIO model, constructed upon causal chains, together with the DIEGO method that strategically rectifies key segments of those chains, offers a viable way to overcome this limitation.
@@ IX Solving Generative RPM Problems
Generative RPM problems represent an advanced variant of RPM tasks. In these problems, the solving models are required to generate at least one solution that conforms to the progressive patterns within the given context matrix, rather than selecting from a predefined bounded option pool, thereby more accurately reflecting the model’s comprehension of the reasoning logic. As solution identification remains progressive-pattern-driven, generative RPM problems inherit the multi-solution characteristic from conventional RPM problems. Consequently, models are expected to produce multiple solutions that are structurally diverse, which poses significant challenges for most traditional neural networks.
@@ IX-A WORLD Guides DIO on Generative RPM problems
Beyond enabling DIO to solve discriminative RPM tasks, the WORLD method also equips it with the capability to tackle generative RPM problems. This method constructs a complete generation pipeline for DIO: “estimating solution distribution $\rightarrow$ sampling solution distribution $\rightarrow$ decoding samples”.
The WORLD method extends DIO by incorporating an external GMM $\{U_{m}|m\in[1,M]\}$, an additional loss term $\ell_{\text{Weighted}}$, and a decoder $\mathcal{D}(\cdot)$. Once trained in this framework, DIO can not only discriminate among the options for each context matrix but also directly generate the corresponding solutions. Specifically, as shown in Formula (42), the loss term $\ell_{\text{Weighted}}$ prompts DIO to assign a context-dependent weight to each GMM component $U_{m}$, where the weight reflects the likelihood that $U_{m}$ contains the correct options for the given context matrix. Within this loss term, the weight for each $U_{m}$ given the context matrix features $\{z_{ij}|i\in[1,8]\}$ is calculated as:
@@
Since these weights $\{S_{mj}|m\in[1,M],j\in[1,N]\}$ indicate the likelihood that the components cover the correct options, we not only gain the chance to sample diverse incorrect options from low-weight components to tighten the lower bound on the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$, but also have the opportunity to combine high-weight components into a solution distribution. To this end, we adopt the Gumbel-Max trick [36] to sample $N$ high-weight components as the solution distribution that DIO estimates for the context matrix $\{x_{i}|i\in[1,8]\}$. Specifically, this trick is applied as follows:
@@
$\text{where}\quad\epsilon_{ij}\sim N(0,1)$
@@
In a nutshell, WORLD requires DIO to interact with a GMM that fully encodes RPM image features, a demand that far surpasses traditional discriminative RPM tasks, thereby sparking DIO’s capacity to solve the generative RPM problems.
@@
Notably, WORLD is well-aligned with the multi-solution characteristic of generative RPM problems. Formulas (45)–(47) indicate that the Gumbel-Max trick ensures the estimated solution distribution $\{U_{\hat{m}_{j}}|j\in[1,N]\}$ exhibits diversity. Formula (48) further reveals that the solutions decoded from samples $\{U_{\hat{m}_{j}}+\epsilon_{ij}|j\in[1,N]\}$ of the estimated distribution have the potential to incorporate the sampling diversity $\epsilon_{ij}$. By iterating the “estimating $\rightarrow$ sampling $\rightarrow$ decoding” pipeline, which can generate diverse solutions, WORLD systematically explores multiple potential solutions and fully embraces the multi-solution logic of generative RPM tasks.
IX-B Evaluating the Generated Solutions
Verifying the solution accuracy of generative RPM problems has long been challenging. Generative RPM problems possess the multi-solution characteristic, which means that the generative solving models cannot be evaluated simply by comparing each generated solution with the provided correct option. This also makes it challenging for RPM experts to devise a perfect verification algorithm. Most prior work on generative RPM problems has used self-verification, i.e., using their own discriminative models to evaluate their own generative models. However, this paper argues that this self-verification paradigm is inadequate. In generative RPM problems, the multi-solution characteristic stems from the progressive-pattern-driven solution evaluation approach. This implies that the algorithms or models used to validate the accuracy of solutions for such problems should also be progressive-pattern-driven rather than mutual-information-driven. The DIEGO approach meets this expectation, as the DIO model trained solely with the loss term $\ell_{\text{DIEGO}}$ will be a progressive-pattern-driven solving model.
More importantly, WORLD is a strong mutual-information-driven inductive model, yet it cannot escape the inherent drawback of neglecting reasoning logic. Using the progressive-pattern-driven model DIEGO to verify WORLD’s ability to solve generative RPM problems not only avoids the imprecision of symmetric verification but also examines the extent to which WORLD’s performance gains are merely superficial.
@@ X Experiment
All experiments conducted in this paper were programmed in Python based on the PyTorch[37] framework. This paper proposes four distinct configurations for RPM problems:
@@ X-3 WORLD-augmented DIO
In the WORLD framework, the hyperparameter $M$ is set to 512, equipping the GMM employed for DIO with 512 components. This configuration enables the generation of $M^{N}=512^{16}$ distinct solution distributions, theoretically allowing for the sampling of an infinite diversity of incorrect options. It is worth noting that when WORLD is used to assist DIO in solving generative RPM problems, $M$ is set to 2048;
@@ X-4 DIEGO-guided DIO
This paper conducted experiments on the RAVEN [18], I-RAVEN[32] and PGM [19] datasets, centered around these four configurations. To ensure fair comparison, we uniformly align the optimizer hyperparameters, dataset scales, and batch sizes for optimizing the aforementioned four configurations with those adopted by the current state-of-the-art RS-Tran model [29].
@@ X-A Experiment on RAVEN and I-RAVEN Problems
The performance of our four configurations alongside previous models on the RAVEN and I-RAVEN datasets is summarized in Table I.
@@
Model Average Center 2 $\times$ 2 Grid 3 $\times$ 3 Grid L-R U-D O-IC O-IG
@@
Triple-CFN[31] 98.9/99.1 100.0/100.0 99.7/99.8 96.2/97.5 99.8/99.9 99.8/99.9 99.9/99.9 97.0/97.3
DIO 98.8/99.3 100.0/100.0 99.6/99.7 96.1/97.4 99.8/99.9 99.8/99.9 99.8/99.8 96.7/97.9
DIO+Brando 99.4/99.5 100.0/100.0 99.8/99.8 98.4/99.0 100.0/100.0 99.9/100.0 100.0/100.0 98.6/98.9
DIO+WORLD 99.5/99.7 100.0/100.0 99.8/99.8 98.8/99.5 99.9/100.0 100.0/100.0 100.0/100.0 99.0/99.5
@@
These results indicate that all four model configurations proposed in this study outperform existing baselines on both the RAVEN and I-RAVEN datasets. Notably, the DIO+DIEGO configuration achieves a peak accuracy of 99.9%, delivering significant performance gains over prior state-of-the-art models (Triple-CFN[31] and RS-TRAN[29]). Particularly in complex reasoning tasks such as $\times3$ Grid and O-IG, the DIO series models demonstrate distinct advantages, validating the generalizability and robustness of their architectural design for abstract visual reasoning challenges.
More critically, the performance achieved by the DIEGO method has overturned a key experimental conclusion in prior studies[31, 23, 29], which posited that the progressive pattern supervision directly implemented from metadata would inevitably undermine the model’s reasoning capabilities. This logically flawed assertion is obsolete. Furthermore, our experiments demonstrate that when no additional supervision signals are incorporated, the strength of mutual information inferred by the solving model serves as the primary determinant of its problem-solving performance, offering critical architectural insights for future abstract reasoning systems. When progressive pattern descriptions are employed as supervisory signals, the supervision utilization strategy combining causal chain modeling with correction of key chain segments proves highly effective.

### v7

@@ VIII-A Target Reasoning Rather Than Solving Problems
This paper realizes that setting the maximization of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective of DIO is suboptimal. Although this statistical measure strongly correlates with the solving accuracy of RPM tasks, it neglects the human-predefined reasoning logic [33], yielding a model optimized for high accuracy rather than faithful causal reasoning. Thus, the performance gains brought by mutual information may be superficial. Both the Brando and WORLD methodologies aim to assist DIO in maximizing mutual information, yet neither resolves this limitation. In a nutshell, mutual-information-driven approaches are effective but need an explicit causal perspective.
@@
In order to verify semantic alignment, this paper takes the $\times3$ Grid” sub-task from the RPM-style RAVEN benchmark [15] as the research object. We conduct training on this dataset using DIO, as well as DIO supplemented with Brando and WORLD. Subsequently, we extract 30 instances that share strictly identical progressive patterns and compute the pairwise similarities among the progressive-pattern representations that DIO encodes for these 30 instances. The calculated $30\times 30$ similarity matrices are visualized as heatmaps in Figure 9.
Figure 9: Heatmap of Progressive-Pattern Representation Extracted from the $\times3$ Grid under RAVEN Problems.
@@
In preliminary experiments, we found that DIO exhibits deviations in modeling the segment of the causal chain that involves “abstract attributes $\rightarrow$ progressive patterns of attributes”. This corresponds to a misalignment between the outputs $P_{\alpha{jkl}}$ produced by Formula (IV-C) and human semantics. Therefore, this paper proposes rectifying the progressive-pattern representations $P_{\alpha{jkl}}$ encoded by DIO as an effective remedy.
@@
To this end, we propose the DIEGO method. DIEGO constructs a set of trainable reference vectors $\{V_{f}|f\in[1,F]\}$. By assigning each vector a distinct metadata modality, this entire set can be considered a comprehensive metadata dictionary. An auxiliary loss term $\ell_{\text{DIEGO}}$ is introduced to align every progressive pattern representation $P_{\alpha jkl}$ extracted by DIO with its corresponding reference vector in $\{V_{f}\mid f\in[1,F]\}$ according to the accompanying metadata. This term is formulated as follows:
@@
Here, $\text{Meta}_{\hat{f}l}$ is a binary indicator that flags whether the $l$-th metadata entry matches the modality associated with the $\hat{f}-$th vector within the reference dictionary $\{V_{f}|f\in[1,F]\}$. Meanwhile, $\tau$ is a trainable scaling coefficient. The detailed formulation of $\ell_{\text{DIEGO}}$ elucidates our alignment mechanism. Guided by the metadata-derived indicators, this loss term employs logarithmic smoothing and exponential scaling to regulate the cosine similarities between each $P_{\alpha_{jkl}}$ and all vectors in the reference dictionary.
Evidently, our objective has shifted from statistical quantity control to vector similarity control, distinguishing DIEGO from Brando and WORLD in that it is no longer mutual-information-driven but progressive-pattern-driven. When the DIO model is trained with the composite loss $\ell_{\text{DIO}}+\ell_{\text{DIEGO}}$ and the trainable reference dictionary $\{V_{f}|f\in[1,F]\}$, the forward pass of the relevant modules proceeds as illustrated in Figure 10.
@@ IX Solving Generative RPM Problems
Generative RPM problems represent an advanced variant of RPM tasks. In these problems, the solving models are required to generate at least one solution that conforms to the progressive patterns within the given context matrix, rather than selecting from a predefined and bounded option pool, thereby more accurately reflecting the model’s comprehension of the reasoning logic. As solution identification remains progressive-pattern-driven, generative RPM problems inherit the multi-solution characteristic from conventional RPM tasks. Consequently, solving models are expected to produce multiple solutions that are structurally diverse, which poses significant challenges for most traditional neural networks. However, WORLD endows DIO with the capability to effectively tackle generative RPM problems.
@@ IX-A WORLD Guides DIO on Generative RPM problems
Beyond enabling DIO to solve discriminative RPM tasks, the WORLD method also equips it with the capability to tackle generative RPM problems. This method constructs a complete generation pipeline for DIO: “estimating the solution distribution $\rightarrow$ sampling the solution distribution $\rightarrow$ decoding the sampling results”.
The WORLD method extends DIO by incorporating an external GMM $\{U_{m}|m\in[1,M]\}$, an additional loss term $\ell_{\text{Weighted}}$, and a decoder $\mathcal{D}(\cdot)$. Once trained in this framework, DIO can not only discriminate among the options for each context matrix but also generate the corresponding solutions. Specifically, as shown in Formula (VII-C), the loss term $\ell_{\text{Weighted}}$ prompts DIO to assign a context-dependent weight to each GMM component $U_{m}$, where the weight reflects the likelihood that $U_{m}$ contains the correct options for the given context matrix. Within this loss term, the weight for each $U_{m}$ given the context matrix features $\{z_{ij}|i\in[1,8]\}$ is calculated as:
@@
Since these weights $\{S_{mj}|m\in[1,M],j\in[1,N]\}$ indicate the likelihood that the components cover the correct options, we not only gain the chance to sample diverse incorrect options from low-weight components to tighten the lower bound on the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$, but also have the opportunity to combine high-weight components into a solution distribution to provide promising options. To this end, we adopt the Gumbel-Max trick [34] to sample $N$ high-weight components as the solution distribution that DIO estimates for the context matrix $\{x_{i}|i\in[1,8]\}$. Specifically, this trick is applied as follows:
@@
$\text{where}\quad\epsilon_{ij}\sim\mathcal{N}(0,1)$
@@
In a nutshell, WORLD requires DIO to interact with a GMM that fully encodes RPM image features, a demand that far surpasses traditional discriminative RPM tasks, thereby sparking DIO’s capacity to solve generative RPM problems.
@@
Commendably, WORLD matches the multi-solution characteristic of generative RPM problems. Equations (44)–(46) show that the Gumbel-Max trick transfers the diversity inherent in the Gumbel noise $G_{mj}$ to the solution index $m_{j}$, so that the estimated distribution $\{U_{\hat{m}_{j}}|j\in[1,N]\}$ ultimately inherits this diversity. Equation (47) then demonstrates that decoding the samples $\{U_{\hat{m}_{j}}+\epsilon_{ij}|j\in[1,N]\}$ has the potential to embed sampling diversity $\epsilon_{ij}$ directly into the resulting solutions $\hat{x}$. Iterating the estimate $\rightarrow$ sample $\rightarrow$ decode pipeline, WORLD systematically explores multiple valid solutions and fully embraces the multi-solution logic of generative RPM tasks.
IX-B Hierarchical Interaction and EMA-based Optimization.
WORLD minimizes the loss function $\ell_{\text{WORLD}}$ to embed the RPM image features $u_{ij}+b_{ij}$ encoded by DIO into its trainable GMM $\{U_{m}|m\in[1,M]\}$. Evidently, the component count $M$ governs WORLD’s effectiveness: an insufficient $M$ forces DIO to compress intricate image features into a low-capacity GMM, diluting the expressivity of these features, degrading DIO’s performance, and obscuring WORLD’s true potential. Although increasing $M$ alleviates these issues, it also incurs computational overhead that scales linearly with $M$ and quickly drives up the cost of optimizing $\ell_{\text{WORLD}}$. This places the setting of $M$ in a quandary when WORLD confronts demanding generative RPM problems.
To this end, we propose two dedicated improvements for the two loss terms $\ell_{1}$ (i.e., Formula (34)) and $\ell_{3}$ (i.e., Formula (36)) within $\ell_{\text{WORLD}}$: hierarchical interaction and exponential-moving-average (EMA)-based component optimization. Specifically, we adopt a G-head hierarchical interaction mechanism that splits the image feature $u_{ij}$ into $G$ independent heads $\{u^{(g)}_{ij}|g\in[1,G]\}$, partitions the $M$ trainable vectors $\{U_{m}|m\in[1,M]\}$ into $G$ groups $\{U^{(g)}_{m}|m\in[1,M],g\in[1,G]\}$, and then performs head-wise hierarchical optimization with the rewritten term $\ell_{1}$ given by:
$\ell_{1}=\frac{1}{16N}\sum_{i=1}^{16}\sum_{j=1}^{N}\Bigl\|u_{ij}-\operatorname{concat}\bigl[U^{(g)}_{\hat{m}_{g}}\bigr]_{g=1}^{G}\Bigr\|_{2}^{2}$ (48)
$\quad\quad\text{where}\quad\hat{m}_{g}=\underset{m\in[1,M]}{\text{argmin}}\bigl\|u^{(g)}_{ij}-U^{(g)}_{m}\bigr\|_{2}^{2}.$
This enables the same $M$ trainable vectors to represent a GMM with $M^{G}$ components, whereby the capacity grows exponentially in $G$ instead of linearly in $M$, eliminating the need to drastically increase $M$ for greater expressivity. However, symmetrically extending the loss term $\ell_{3}$ in the same fashion would significantly increase optimization overhead. Therefore, rather than supervising the hierarchical training of components $\{U^{(g)}_{m}|m\in[1,M],g\in[1,G]\}$ with $\ell_{3}$, we switch to an EMA-based update strategy that enables these GMM components to evolve without the gradient back-propagation overhead. The EMA-based update for $U_{m}^{(g)}$ is given by:
$\hat{m}_{ij}^{(g)}$ $=\underset{m\in[1,M]}{\mathrm{argmin}}\|u_{ij}^{(g)}-U_{m}^{(g)}\|_{2}^{2}$ (49)
$n_{m}^{(g)}$ $=\sum_{i,j}\mathcal{I}[m=\hat{m}_{ij}^{(g)}],s_{m}^{(g)}=\sum_{i,j}\mathcal{I}[m=\hat{m}_{ij}^{(g)}]\frac{u_{ij}^{(g)}}{n_{m}^{(g)}}$ (50)
$N_{m}^{(g)}$ $\leftarrow\eta N_{m}^{(g)}+\gamma\,n_{m}^{(g)},\,S_{m}^{(g)}\leftarrow\eta S_{m}^{(g)}+\gamma\,s_{m}^{(g)}$ (51)
$U_{m}^{(g)}$ $\leftarrow\dfrac{S_{m}^{(g)}}{N_{m}^{(g)}+\varepsilon}\Bigg/\sum_{k=1}^{M}\dfrac{N_{k}^{(g)}}{N_{k}^{(g)}+\varepsilon}$ (52)
Here, $\eta$ is 0.99; $\gamma$ is 0.02; $\varepsilon$ is $10^{-5}$; $N_{m}^{(g)}$ and $S_{m}^{(g)}$ are initialized to 0; $\mathcal{I}[\cdot]$ denotes the indicator function.
Equations (49) and (50) compute the count $n_{m}^{(g)}$ and the mean $s_{m}^{(g)}$ of the representations $u_{ij}^{(g)}$ assigned to the GMM component $U_{m}^{(g)}$, while Equations (51) and (52) perform the momentum updates of $n_{m}^{(g)}$, $s_{m}^{(g)}$, and $U_{m}^{(g)}$. These updates are executed per gradient step on the $\ell_{1}$ loss term, ensuring that the GMM evolves in lockstep with DIO.
IX-C EMA-Based Efficient Estimation of Solution Distribution
The G-head hierarchical interaction mechanism enables $M$ vectors to support a GMM with $M^{G}$ components; however, as indicated by Equations (43)–(46), it causes a sharp increase in the number of times $\tilde{\text{DIO}}$ must weigh the components, imposing a heavy computational burden.
To this end, a EMA-Based approach is adopted: cluster the $M^{G}$ components into $M$ centroids, assign weights to these centroids using the procedure in Equation (43), and perform Gumbel-MAX sampling on the $M$ weights as indicated by Equation (46); then iteratively repeat clustering, weight assignment, and sampling within the selected cluster until only $M$ components remain. Finally, use Gumbel-MAX to select the component $U_{\hat{m}}$ that can serve as the solution distribution from these $M$ components. This reduces the number of weighting operations that $\tilde{\text{DIO}}$ must perform from $M^{G}$ to $M\times G$.
This recursive scheme demands that $\tilde{\text{DIO}}$ be Lipschitz continuous (i.e., it must assign similar weights to similar components). We therefore apply spectral normalization [35] to every parameter matrix in $\tilde{\text{DIO}}$ and apply Lipschitz normalization [36] to the self-attention mechanism within each block, ensuring the required Lipschitz property.
IX-D Validation of Generative RPM problems
Notably, the validation of generative RPM has long been hindered by its multi-solution nature: any answer that satisfies the progressive pattern can be correct, rendering traditional point-to-point comparisons ineffective. Validation must therefore adopt the same progressive-pattern-driven principle, one that is already embedded in the DIO model trained with the $\ell_{\text{DIEGO}}$ loss, enabling it to perform verification effectively.
@@ X Experiment
All experiments in this paper were conducted on four A100-80G GPUs using Python and the PyTorch[37] framework. This paper proposes four configurations for RPM problems:
@@ X-3 WORLD-augmented DIO
In the WORLD framework, when solving discriminative RPM problems, $M$ is set to $2^{11}$, and a random window of size 256 is defined to select the components used for computing the loss term $\ell_{\text{Weighted}}$. If the $G$-head hierarchical interaction is activated, $G$ is set to 2, $M$ remains $2^{11}$, and each head employs a window of size 16 to compute the $\ell_{\text{Weighted}}$ loss. When the framework is applied to generative RPM tasks, $G$ is set to 2, and $M$ is set to $2^{14}$; the random window for $\ell_{\text{Weighted}}$ inside each head is kept at 16.
@@ X-4 DIEGO-guided DIO
This paper conducts experiments on the RAVEN [15], I-RAVEN [28], and PGM [16] datasets, centered around these four configurations. To ensure a fair comparison, we use the same optimizer hyperparameters, dataset scale, and batch size as those of the current SOTA RS-Tran model [26].
@@ X-A Experiment on RAVEN and I-RAVEN Problems
Table I summarizes the performance of our four configurations in comparison with previous models on the RAVEN and I-RAVEN datasets, where “WORLDG” denotes the activation of the $G$-head hierarchical interaction.
@@
Model Average Center $\times2$ Grid $\times3$ Grid L-R U-D O-IC O-IG
@@
DIO 98.8/99.2 100.0/100.0 99.6/99.7 96.3/97.4 99.8/99.9 99.8/99.9 99.8/99.8 96.7/97.9
DIO+Brando 99.2/99.5 100.0/100.0 99.8/99.8 97.4/98.6 99.9/99.9 99.9/100.0 100.0/100.0 97.8/98.8
DIO+WORLD 99.4/99.7 100.0/100.0 99.8/99.8 98.5/99.2 99.9/100.0 100.0/100.0 100.0/100.0 98.4/99.3
DIO+WORLDG 99.6/99.7 100.0/100.0 99.8/99.8 99.2/99.5 100.0/100.0 100.0/100.0 100.0/100.0 99.3/99.5
@@
These results indicate that all four configurations outperform the existing baselines on both the RAVEN and I-RAVEN datasets. Notably, the DIO+DIEGO configuration achieves a peak accuracy of 99.9%, delivering significant performance gains over the prior SOTA model (RS-TRAN[26]). Particularly in complex reasoning tasks such as $\times3$ Grid and O-IG, the DIO series models demonstrate distinct advantages, validating the generalizability and robustness of their architectural design for abstract visual reasoning challenges.
Moreover, our experiments demonstrate that when no additional causal supervisory signals are introduced, the strength of the mutual information inferred by the reasoning model becomes the primary determinant of its problem-solving performance, offering crucial architectural insights for future abstract-reasoning systems. When progressive pattern descriptions are employed as supervisory signals, a supervision strategy that combines causal-chain modeling with the correction of key chain segments proves to be highly effective. This overturns the long-standing conclusion of prior studies [20, 23, 22, 26] that direct progressive-pattern supervision inevitably undermines model performance.